OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −1 −4 4 −3 3

OpenStudy (anonymous):

helppp pleasee

OpenStudy (pfenn1):

\[3x + 2y = 6\]\[2x + 3y = −1 \]Solve one of the equations for y in terms of x ans substitute it into the other equation and solve for x.

OpenStudy (pfenn1):

*and

OpenStudy (pfenn1):

2y=6-3x\[y=\frac {6-3x}{2}\]

OpenStudy (pfenn1):

\[2x+3\left( \frac{6-3x}{2} \right)=-1\]Can you solve it from here?

OpenStudy (anonymous):

i dont get this i'm supostu to do this by eliminetion method.

OpenStudy (pfenn1):

This is the elimination method. We "eliminated" the y from the second equation by substituting the expression for y we found from the first equation. I took the first equation and subtracted 3x from both sides, then divided both sides by 2.

OpenStudy (precal):